Home‎ > ‎Crew 153‎ > ‎

414-CommandersReport

Commander's Report
Date: April 14, 2015
Crew Commander Name: Bastien MATHURIN

Hello Earth,

This day on Mars was a good day. Today cooking is in charge of Romain and Martin. They're making amazing pizzas with chorizo, tomato, pepperoni, onions, pineapple and cheese. As crazy as it may seem, it looks really tasty. I hope it will be a great time for us.

We had requested 2 EVAs for today. One early am with a view to take some rock samples in old streambeds for Florian microbiology experience and the other one at 4:00 pm to calibrate Auriane's susceptibilimeter that is needed for her experience about magnetic properties of rocks. The fact is that we encountered some issues during these EVAs:

        -Florian has broken an EVA helmet while using the ATVs. Although they were riding safely on the road, the wind tore the rope maintaining the helmet to the ATV (he was wearing the required protective helmet) and it hit his ATV's wheel. There is therefore only 3 EVA helmets available. For the future, we will try to find a new solution to avoid this problem.
        -The second EVA had to be shortened due to the wind that was too strong. Auriane just had the time to calibrate her tool. She would have liked to make some preliminary measurements but it was impossible with that wind.

I told in yesterday CapCom that our Crew Journalist Romain didn't feel very well. He feels a bit better today but his health isn't perfect at all. He takes some medicines to solve this issue but there won't be drugs at all in a few days for his pain... All the team hopes that he will feel better in a few days otherwise we will maybe have to break the sim to see a doctor for him.

Our Crew Geologist, Auriane, worked on the backpacks that were not working well. The battery of backpack 1 was not working perfectly and she replaced it. On backpack 4, she completely changed the fans and the straps. All backpacks should now be operational for future EVAs.

Concerning the technical issues we had previously, there is no change for the water supply in the Hab. We're still working with our crew engineer temporary solution. I hope that a new pump will be delivered as soon as possible (we don't have water in the ground floor, so it not convenient at all for the restrooms and the shower). However, the issue about the telescope is solved since yesterday night: it was the remote control that was defective and our Crew Astronomer, Martin, just had to replace it. He should therefore be able to observe the stars this night.

Despite all the problems listed hereinabove, our crew is having a great time on Mars!

Ad astra!